Success on the Uno!

Author: Lars Larson

The CO2 sensor problems from earlier this week were due to a faulty breadboard which left insufficient current for the sensor.  After swapping the whole array to a new board, the CO2, temp, RH, and SD card are up and running.

Average values in the OPEnS Lab today:

CO2: 835 ppm, temp: 23˚C, RH: 45%

Next steps: coordinating with Elad to set up a more rigorous testing protocol, and wiring in the luminosity sensor when it arrives.